The ice and steam point of a mecury in gas thermometer of centigrate scale &of uniform bore corresponds respectively to 3cm &19cm length of the mecury thread when the length is 12cm what will the temperature be?
1
Expert's answer
2018-02-14T08:49:06-0500
We can use the formula for calculating the temperature on an unmarked thermometer:
θ=(x-x_0)/(x_100-x_0 ) 100℃
In our case:
θ=(12-3)/(19-3) 100℃=56.25℃